Can someone please help me solve this linear system in three variables?
x-3y=11
5y+z=4
-x+z=2
{(?,?,?)}

Answer by scott8148(5892)

adding 1st and 3rd eqns __ z-3y=13 __ adding 3y __ z=3y+13

substituting into 2nd eqn __ 5y+(3y+13)=4 __ subtracting 13 __ 8y=-9 __ y=-9/8

substituting into 1st eqn __ x-3(-9/8)=11 __ subtracting 27/8 __ x=61/8

substituting into 3rd eqn __ -(61/8)+z=2 __ adding 61/8 __ z=77/8